OpenStudy (anonymous):

please help!!! need now !!! The data gives the responses of 15 students who were asked how many text messages they send in a day. (2, 50, 50, 51, 51, 52, 53, 54, 54, 55, 55, 55, 56, 56, 57} What is the best measure of the central tendency of the data? mean mode median range interquartile range

OpenStudy (anonymous):

mean = 704(sum of values)/15(number of values) = 46.933 mode = value that appears most = 55 median = the middle value = 54 range = the difference between the highest and lowest values = 55 interquartile range = long explanation, go here for more info www.statistichowto.com/interquartile-range/ = 55 -51 = 4

OpenStudy (anonymous):

the median